OpenStudy (anonymous):

can someone help me with this problem dx/dt = sqrt(1+sin(t^2)). I only know how to separate variables sqrt(1+sin(t^2))dt = dx

OpenStudy (anonymous):

The problem statement, all variables and given/known data Find the area of the surface obtained by rotating the curve about the x-axis: y=cos 2x, 0<=x<=pi/6 2. Relevant equations Surface area about the x-axis = Integral of 2pi * f(x) * sqrt(1+[f'(x)]^2) dx 3. The attempt at a solution I think I set up my integral correctly, so my problem is figuring out how to solve the integral. I have: 2pi * Integral of cos 2x * sqrt(1+4sin^2 2x) dx on [0, pi/6] I can't figure out how to solve that. And can someone teach me how to use Latex for integrals so it looks better.

OpenStudy (anonymous):

I haven't learned how to find surface about the x-axis. I'm just trying to differentiate

sam (.sam.):

separate then differentiate

sam (.sam.):

dx/dt = sqrt(1+sin(t^2)) dx= sqrt(1+sin(t^2)) dt

sam (.sam.):

C(x) is fresnel C integral and S(x) is fresnel S integral

OpenStudy (anonymous):

is there any way to do it with the Fundamental Theorem?